# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v2

@@ B.2 Proof of Lemma 5
In this section, we consider how to measure the difference of ${\mathbf{x}}$. To be more specific, we mainly consider the relationships between the following 44 44 terms:
$\Delta_{1}^{(k)},\Delta_{2}^{(k)},\mathbb{E}_{k}[\|\Delta_{x}^{(k)}\|_{F}^{2}],\Delta_{y}^{(k)},\quad k\in[m^{*}].$
Lemma 7 will give upper bounds of $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ using other terms, while Lemma
Lemma 7.
Suppose Proposition 2 and Assumption 34 hold. When $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold for any positive integers $k$ and $m$.
@@
Proof.
Define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1}$, when $j\geq 1$. We have
@@
$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+ms_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
@@
where in the first inequality, the $\sigma^{2}$ terms are separated using Lemma 3, and the second inequality comes from rolling sum.
$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}$
$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}$
For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$. This gives
$\displaystyle\quad\Delta_{2}^{(k)}\leq 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
@@
$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (34)
@@
Plug (34) into the estimate for $\Delta_{1}^{(k)}$, it is not difficult to verify that $\Delta_{1}^{(k)}$ could use the same upper bound as (34).
Next, we provide some inequalities for the consensus error $\Delta_{x}^{(k)}$ and $\Delta_{y}^{(k)}$, $k\in[m^{*}]$.
For $k\in[m^{*}]$, we have
@@
Define
$\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)}).$
Note that it can be rewritten as
@@
according to Lemma 3, we have
$\displaystyle\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$
where
@@
By applying the rolling sum lemma, we have
@@
Plug in (33), we have
$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}$
$\displaystyle\,\,+6s_{B^{m}}nL\Delta+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
@@
Define
$\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)},$
then we can iterate (38) and obtain that
$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(k)}.$
To further analyze $\Delta_{4}^{(k)}$, we expand it as
@@
$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
@@
$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]$
$\displaystyle\quad+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
Using the rolling sum lemma, we have
@@
$\displaystyle\frac{c\alpha L^{2}}{2}\quad\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq\frac{\alpha cm}{4}C_{\sigma,1}K\sigma^{2}+\frac{c\alpha}{4}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{c\alpha L\Delta}{4m},$
@@
$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
This finishes the proof.
@@ B.3 Proof of Theorem 1
$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)$
@@ C.1 Weight Matrix Design
We use the following procedure to generate a row/column-stochastic weight matrix. For a fixed number of nodes $n$, we begin by initializing a 0-1 adjacency matrix $W$ as the identity matrix $I$ to account for self-loops on each node. Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$. After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from 11 11 to 99 99 to all entries where $W_{i,j}=1$. Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
@@ C.2 Synthetic data
For experiments conducted on the synthetic dataset, the objective function is given by:
@@
Data Synthesis Process:
  1. 1.
    Global Data Generation:
    • Generate optimal parameters: $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$
    • Create global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with $h_{l,j}\sim\mathcal{N}(0,1)$
    • Compute labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ via randomized thresholding:
@@
  2. 2.
    Data Distribution:
    • Partition $H$ and $L$ across $n$ nodes using:
      $H^{(i)}=H[iM:(i+1)M,\;:\;],\;\quad Y^{(i)}=Y[iM:(i+1)M],\;\quad M=L_{\text{total}}/n$
      where $L_{\text{total}}$ must be divisible by $n$
  3. 3.
    Local Initial Points:
    $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d}),\quad\sigma_{h}=10$
Gradient Computation: At each iteration, each node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$. The gradient computation consists of two components:
@@
where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly from the $L_{i}$ local samples without replacement. Notice that the gradient on each node $i$ is computed on the local parameter $x_{i}\in\mathbb{R}^{d}$.
Implementation Details:
  • Global dataset size $L_{\text{total}}=204800$, batch size $B=200$
  • Dimension $d=10$, regularization $\rho=0.01$
  • Node configurations: $n\in\{4,6,8,9,12,16,18,24,25\}$, and the least common multiple (LCM) is $3,600$.
Notice that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
@@
Evaluation Metric: We track $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ as gradient norm, where
@@
$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\;M=L_{\text{Total}}/n.$
The experimental results shown in our plots represent the average performance across 2020 2020 independent repetitions, where each trial is conducted with the fixed random $\text{seed}=42$ for reproducibility.
C.3 Neural Network Experiment
We first employ three-layer fully connected neural networks for MNIST classification.
Data Distribution: For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector
$\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$
where $\mathbf{1}^{10}$ is a 1010 1010-dimensional vector of ones. Each $q_{k,c}$ represents the target proportion of samples from class $c$ that node $k$ should ideally hold. By definition of the Dirichlet distribution, $q_{k,c}\geq 0,\forall c$ and $\sum_{c=0}^{9}q_{k,c}=1,\forall k$.
Denote $N_{k,c}^{\text{target}}=q_{k,c}\cdot\frac{M}{n}$ as the target number of samples of class $c$ to be assigned to node $k$. Ignoring the slight deviations caused by non-divisibility, the number of samples assigned to each node $k$ is approximately $60,000/n$. Here we set $\alpha=0.9$ to get the heterogeneous data and the Figure 5 shows the distribution of samples with 1616 1616 nodes. The Figure 2 shows that the Push Pull achieves linear speedup on synthetic data.

### v3

@@ B.2 Proof of Lemma 4
We first provide the complete definition of the constant $C_{m}$ used in the statement. We define $C_{m}$ as the maximum of all lower bounds imposed on $m$ to ensure the validity of subsequent inequalities and simplifications:
$\displaystyle C_{m}=\max\Big\{$ $\displaystyle n^{-1}s_{A}^{2},20c^{-2}ns_{A}s_{A^{m}},\;c^{-2}n^{2}s_{B}^{2},30c^{-2}s_{B}^{2}s_{B^{m}}$
$\displaystyle c^{-2}n^{2}(8s_{A}^{2}s_{B}+2ns_{B}+2s_{A}s_{B}+\sqrt{n}s_{B}),20c^{-2}n^{3}s_{A}s_{A^{m}}s_{B}s_{B^{m}}\Big\}.$ (42)
We define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1},j\geq 1$ and the following constants:
$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}\leq 8s_{A}^{2}s_{B}+2ns_{B},$
$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}\leq 2s_{A}s_{B}+\sqrt{n}s_{B}.$
For all arguments used in this section, we suppose Proposition 2 and Assumption 34 hold. We first declare that, when $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),n^{-1}s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold $\forall k,m\geq 1$.
@@
Note that
@@
$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+mns_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
@@
where in the first inequality, the $\sigma^{2}$ terms are separated using Lemma 6, and the second inequality comes from rolling sum. For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$. This gives
$\displaystyle\quad\Delta_{2}^{(k)}\leq$ $\displaystyle 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}$
$\displaystyle+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
@@
$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (45)
@@
Plug (45) into the estimate for $\Delta_{1}^{(k)}$, it is not difficult to verify that $\Delta_{1}^{(k)}$ could use the same upper bound as (45). Next, we provide some inequalities for the consensus error $\Delta_{x}^{(k)}$ and $\Delta_{y}^{(k)}$, $k\in[m^{*}]$. For $k\in[m^{*}]$, we have
@@
We define $\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$, which can be written as
@@
according to Lemma 6, we have $\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$, where
@@
By applying the Jensen’s inequality, we have
@@
Replace the $\Delta_{2}$ term with the estimate in (44), we reach that
$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq$ $\displaystyle(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}+6s_{B^{m}}nL\Delta$
$\displaystyle+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
@@
For $k\in[m^{*}]$, define $\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}$, then we can iterate (49):
$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(jm)},\quad k\in[m^{*}].$
Using the rolling sum lemma, we know that for $k\in[m^{*}]$,
$\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]].$
To further analyze $\mathbb{E}_{k}[\Delta_{4}^{(k)}]$, $k\in[m^{*}]$, we expand it as
@@
$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
@@
$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
Plug in our estimate on $\Delta_{4}^{(k)}$, we have
@@
$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq C_{\sigma,1}K\sigma^{2}+\frac{1}{m}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{L\Delta}{m^{2}},$
@@
$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$ (53)
This finishes the proof of Lemma 4. Finally, we leave a brief discussion on $s_{B^{m}}$. Recall that
$s_{B^{m}}=\max\{\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}\},$
when $m$ increases, both $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}$ and $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}$ converge to zero exponentially fast, and $\|I-B_{\infty}\|_{2}\geq 1$. Thus, $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}^{2}$ exponentially fast as $m$ grows.
@@ C.1 Weight Matrix Design
We use the following procedure to generate a row/column-stochastic weight matrix. For a fixed number of nodes $n$, we begin by initializing a 0-1 adjacency matrix $W$ as the identity matrix $I$ to account for self-loops on each node. Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$. After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from $1$ to $9$ to all entries where $W_{i,j}=1$. Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
@@ C.2 Synthetic data
For experiments conducted on the synthetic dataset, we minimize the objective:
@@
Data Synthesis Process. We generate the synthetic dataset through the following three-step procedure. (1) Global Data Generation: We first generate optimal parameters $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$ and create a global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with entries $h_{l,j}\sim\mathcal{N}(0,1)$. Labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ are computed via randomized thresholding:
@@
(2) Data Distribution: We partition $H$ and $Y$ equally across $n$ nodes using $H^{(i)}=H[iM:(i+1)M,:]$ and $Y^{(i)}=Y[iM:(i+1)M]$ where $M=L_{\text{total}}/n$, requiring $L_{\text{total}}$ to be divisible by $n$. (3) Local Initial Points: Each node $i$ starts from $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i}$ where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d})$ with $\sigma_{h}=10$.
Gradient Computation. At each iteration, node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$:
@@
where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly without replacement from the $L_{i}$ local samples. The gradient at each node $i$ is computed on its local parameter $x_{i}\in\mathbb{R}^{d}$.
Implementation Details. We use global dataset size $L_{\text{total}}=204800$ with batch size $B=200$, dimension $d=10$, and regularization parameter $\rho=0.01$. Node configurations include $n\in\{4,6,8,9,12,16,18,24,25\}$ with least common multiple (LCM) of $3,600$. Note that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
@@
Evaluation Metric. We track the gradient norm $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ where
@@
$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\quad M=L_{\text{Total}}/n.$
Results represent averages over 20 independent trials with fixed random seed 42 for reproducibility.
C.3 Neural Network Experiments
We evaluate our Push-Pull method on two benchmark tasks: MNIST classification using three-layer fully connected networks and CIFAR-10 classification using ResNet-18.
Data Distribution. To simulate heterogeneous data distributions across nodes, we employ Dirichlet sampling. For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector $\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$, where $\mathbf{1}^{10}$ is a 10-dimensional vector of ones. Each $q_{k,c}$ represents the proportion of samples from class $c$ assigned to node $k$, satisfying $q_{k,c}\geq 0$ and $\sum_{c=0}^{9}q_{k,c}=1$. The target number of samples of class $c$ for node $k$ is $N_{k,c}^{\text{target}}=q_{k,c}\cdot M/n$, resulting in approximately $60,000/n$ samples per node for MNIST and $50,000/n$ for CIFAR-10. We set $\alpha=0.9$ to create heterogeneous distributions, as illustrated in Figure 5 for 16 nodes.
Evaluation Metrics. We measure convergence using the normalized gradient norm computed after each batch. Specifically, we average the gradients from all $n$ nodes and normalize by the square root of the total parameter count:
$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{j}}\right\|_{2}\Big/\sqrt{P},$ (57)
where $\mathbf{g}_{i,B_{j}}$ is the gradient computed on the $j$-th batch of node $i$, and $P$ is the number of parameters. For MNIST with three-layer FCNs, $P=1,863,690$; for CIFAR-10 with ResNet-18, $P=11,173,962$.
Experimental Setup. We test our method on various network topologies including directed exponential and grid graphs, as well as undirected random graphs. Training uses learning rate $lr=0.005$ and batch size $B=128$ for MNIST, $B=128$ for CIFAR-10, consistent across all topologies and node configurations. Performance is compared against single-node baseline training ($n=1$). Figures 2 and 3 demonstrate that Push-Pull achieves linear speedup on both synthetic data and real datasets across directed and undirected topologies. Figure 4 shows the validation accuracy comparison for CIFAR-10.